### v1

# CCoE: A Compact LLM with Collaboration of Experts

Abstract
In the domain of Large Language Model (LLM), LLMs demonstrate significant capabilities in natural language understanding and generation. With the growing needs of applying LLMs on various domains, it is a research question that how to efficiently train and build a model that has expertise in different domains but with a low training cost. We propose CCoE architecture, a framework of easily coupling multiple strong domain experts together to fuse into a big LLM, provides a collective way of utilizing the different domain expert LLMs. Besides, training a large collaborative of multiple expert LLMs requires a high requirements on training sources. CCoE bypasses this problem through isolating other experts and train each expert separately. The design of CCoE assembles multiple expert LLMs through the CoE (Collaboration of Experts) layer. Each CoE layer could have one or more expert LLMs. Expert LLMs have different number of layers and have been well-trained for different domain tasks. Each expert is fine-tuned to be able to achieve the comparable results with SOTA domain LLMs. We start from 5 experts in the domain of Code, Math, Law, text-to-SQL and Medical. The results indicate that our CCoE framework can easily and efficiently boost nearly 10%-20% performance on original base model in different domains but using less resources on training, as well as inference.
1 Introduction
Large Language models (LLMs) have shown great advances in diverse domain tasks, primarily due to they have been trained with vast amount of high quality data and aligned with human preferences. The pre-training and human preference alignment process enable LLMs to demonstrate the extraordinary capabilities in natural language process and problem-solving. Despite the great success, most existing open-sources LLMs are still far way from giving satisfactory results in solving problems like code, mathematics because those domains requires a strong reasoning ability. Another limitations of pre-trained LLMs is the lacking of specialized domain knowledge like Medical due to limited domain corpus in pre-training phrase. To improve LLM’s abilities in mathematical reasoning, code generation or knowledge in specialized domain, these can be further enhanced through supervised fine-tuning.
From another perspective, the current LLMs shows that they tend to demonstrate to have strength and weakness in various tasks and domains based on their evaluations results. some LLMs are excellent in common sense logical reasoning [1]. Other LLMs are well-suited for code generation [2] or mathematical problem solving [3], [4]. However, it is great challenging task for training a model that is well-suited in all different domains. This intrigues to ask the question: Can we boost the model‘s performance in all domains with low training cost?
Recent work focuses on collaborating or assembling multiple LLMs with different sizes to form a more capable and robust model becomes a new existing research direction. Since different LLMs demonstrate the strength and weakness in different tasks and fields, it has great potentials to utilize the different LLMs’s strength in various tasks. [5] [6]. Another framework is to ensemble multiple LLMs’ responses at each decoding step [7] or creating a voting mechanism to select the top answer at the final step. This has already done to use multiple LLMs to generate training dataset [8] or final answer [9]. These recent work indicates the phenomenon of collaborativeness that allows LLMs working together to generate better results. Therefore, it is a rising tendency of creating an unified framework that enables seamlessly multiple LLM experts to collaborate together for open-ended tasks. We propose the CCoE framework that incorporates the idea of effectively coupling multiple expert LLMs together as a fused LLM to greatly improve the overall performance of original base model in all domains. In CCoE framework, we define each expert as several layers of LLM. The expert layers are well-trained for domain-based task to enhance the domain capabilities and could have flexible number of parameters in the network. Similar to MoE, the expert layers form a specialized sub-network and can form a complete LLM with interconnecting to some layers of base model. CCoE defines the fine-grained functionality of each expert to solve different open-ended tasks or being knowledgeable in certain domains. We solve this problem by coupling each expert at CoE layer without any modifications of original model. We propose two types operations called "pop" and "push" to allow each expert to be fine-tuned separately. As for expert routing, we adapts the rule-based gating mechanism, which is discussed in section 3.1. Furthermore, at each CoE layer, experts are loosely coupled, which provides the scalability. CCoE framework supports to flexibly scale up to any amount of experts. Our main contributions are as following:
  • CCoE provides a framework that fuses multiple expert layers into one LLM to boost the performance in all domains. It can easily scale up the number of experts to increase the model capabilities.
  • CCoE does not need to activate all the parameters. It only needs to activate a portion of its parameters to generate answer for each prompt query.
  • CCoE provides the flexibility of supporting experts training and continual training without impacting to the overall model.
  • Our experiments indicate that experts can easily outperform the base model in all domains and even achieve the comparable results with other domain LLMs through efficient fine-tuning on high quality well-designed domain dataset.
2 Analogy to Mixture of Experts
The MoE network usually built based on Transformer architecture that mainly consists of one attention layer, one Feedforward layer and two normalization layers. A standard MoE layers contains multiple Feedfoward layer $E_{1},\cdots,E_{n}$ and each Feedforward layer is an Expert $E_{i}$. Each token is generally assigned to one or two experts. The gating network decides the token assignment for each expert, which is demonstrated in Equation (1).
$y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x)$ (1)
, where $G(x)$ and $E_{i}(x)$ are the outputs of the gating network and the $i$-th expert with the given input $x$. Commonly, $G(x)$ is a sparse vector so that there is not need to compute $E_{i}(x)$ when $G(x)$ has the value of 0. The design of gating network can simply uses a Softmax gating [10]. It multiplies a trainable weight matrix on the input and then applies the $Softmax$ function (Equation 2).
$G(x)=Softmax(x*W_{g})$ (2)
Softmax design of gating can easily introduce the expert unbalancing problem during the training: the routing network always select one or two experts during the training, whereas rest of experts cannot update their weights. Noisy Top-K Gating can mitigate this problem [11] (Equation 3).
$\begin{split}G(x)=Softmax(TopK(H(x)),k))\\ \\ H(x)_{i}=(x\cdot W_{g})_{i}+StdNorm()\cdot Softplus(x\cdot W_{noise}+i)\\ TopK(v,k)_{i}=\begin{cases}v_{i},\quad&v_{i}\in\text{TopK of v}\\ -\infty,\quad&otherwise.\end{cases}\end{split}$ (3)
During the training, the effect of noisy efficiently avoids the problem that the gating network always selects the same expert. From the high-level perspective, our proposed CCoE framework integrates the experts at the model level, whereas the MoE builds the experts at the layer level (Feedfoward layers). With coupling to the base model, each expert can be viewed as a complete and compact LLM that is trained for specific downstream task. Although CCoE and MoE all have the specialized sub-networks within a single model, our CCoE clearly clarifies the functionality of each sub-network. This gives a higher interoperability of understanding the working flow of the model. Moreover, CCoE reduces computational overheads associated with fine-tuning model to enhance one of downstream domain capabilities. It also provides higher scalability through adding more experts to support a new fine-grained downstream tasks and open-ended questions. For example, it can easily add an expert for targeting English-Chinese translation task without worrying about the task conflicts with other experts.
3 Dilemma of Supervised Fine-tuning LLM
Large Language model (LLM) Supervised fine-tuning (SFT) is the process of continually training the pre-trained models on smaller, specific datasets to refine their capabilities and improve the performance in a specific task. SFT can also be used to inject new domain knowledge into the model. The SFT process is about turning a general-purpose model and transform it into specialized model with the domain-related dataset. For example, there is a limited knowledge in pre-trained model on law domain. We can inject the knowledge of the law into LLM to make it become a law specialized model. The SFT brings a gap between pre-trained model and fine-tuned model. Nowadays, LLM fine-tuning has become an indispensable approach for enterprises to enhance their operational process. Through training LLMs for specific task with domain datasets, we can push the knowledge of LLMs to the boundaries of different areas [12] [13] [14].
There are many ad-hoc attempts on SFT for enhancing LLM performance on individual capabilities in open community. However, when the widely used LLMs have been fine-tuned, it may affect the general knowledge stored in LLMs since there is a distribution shift between SFT data and the original training data. The study of SFT is crucial for certain practical applications of LLMs. In most real cases, users want the LLMs to be enhanced in specific domain capabilities while preserving their general capabilities. A significant challenge in this paradigm is catastrophic forgetting (CF), in which a model forgets previously learned knowledge during the SFT process [15] [16]. The naive way of reducing the CF is to mix the supervised data with the pre-trained data at certain ratios. Recently, many researches have proposed various approaches to alleviate the problem of CF such as Dual-Stage Mixed Fine-tuning [17], Recall and Learning [18], etc. However, there are still many remaining challenges left to the community.
4 Collaboration of Expert
Section 4 discusses the methodology of CCoE from the design and implementation details. We first demonstrate the CCoE structure and its variations. We also present that how CCoE solves the dilemma of LLM SFT from the design perspective. Then, we study the objectives of formulating the CCoE model to boost the overall performance of base model and the problem of balancing the performance and resources.
4.1 Design of CCoE
CCoE mainly adopts the idea of Collaboration-of-Experts. In Figure 1, CCoE consists of one backbone LLM and $n$ domain experts and can be formed as a total of n+1 complete LLMs. We define the expert as the sub-network of the LLM. A CoE layer is a layer that is conjunct of experts. Each CoE layer could have different number of experts based on overall design. Demonstrated in Figure 1, the first advantage of CCoE framework is the flexibility in terms of expert size and task. Experts could have different sizes. This indicates that each expert could have different number of parameters to balance the performance and the resources. To be more specifically, LLMs face inherent constraints on model size and training data. Some downstream domains may not have enough tokens to train a large size LLM. A small size LLM may be more suitable in this case. A large LLM may be "a storm in the teacup" for some simple tasks. For example, sentiment analysis task is relatively easy comparing to natural language generation and inference. Many recent researches have proposed to use bert-based model instead of LLMs [19].
In CCoE, each expert is coupled at the CoE layer. Given a $n$-layer backbone LLM, each layer could be the CoE layer and each CoE layer could have any number of experts. This demonstrates the scalability of CCoE framework that each CCoE model is able to scale up to as many as number of experts until meets the constraints of computational resources. The third advantage of LLM is expert isolation. CCoE framework allows each expert to train separately. The training corpus of LLM has the issue of being out-of-date or we want LLMs to continue to learn. This induces the concept of continual pre-train that aims to fine-tune the LLM with new knowledge. However, fine-tuning the entire LLM usually is costly especially for those LLMs that have large sizes. CCoE framework can bypass this by continual training one of the expert LLMs or add more experts instead of training the entire model.
Figure 1: A high level structure demonstration for CCoE model framework
We can denote the CCoE framework in Figure 1 as $E_{b,L},E_{1,l_{1}},E_{2,l_{2}},\cdots,E_{3,l_{3}}$, Assume $E_{b,L}$ is the base LLM with total number of $L$ number of layers. One of expert processes an input text to generates its continuation. Formally, for given an input prompt $x$, the output of CCoE can be expressed as following:
$y=E_{i,l_{i}}(E_{b,L-l_{i}}(x))$ (4)
The first $L-l_{i}$ layers of the base model process the input $x$ and then the $l_{i}$ layers of expert $i$ process the output of base LLM $L-l_{i}$ to generate the final answer. The vocabulary expansion is currently out of the scope but it can be expand to if there is necessary.
There are two variations of the CCoE based on the different design of the experts. The first variation is a special case of the CCoE, which all experts share the same layers of base model, but experts can still have different number of layers. Then, the output of this special case CCoE can be formally written as:
$y=E_{i,l_{i}}(E_{b,l_{b}}(x))$ (5)
Here, $l_{b}$ is the number of shared layers in the base model and $l_{i}$ is the expert layers. The input prompt will always be processed by the first $l_{b}$ layers of the base model and then routes to the expert. The second variation of the CCoE is to refactor the MoE structure. It has almost the same structure as the MoE except the routing network. Expert are still coupled at the Feedforward layer but the routing network is replaced by rule-based gating. The output of CCoE for this variation can be written as:
$h=LayerNorm(X_{in}+Attention(X_{in}))$ (6)
$y=LayerNorm(h+E_{i,l_{i}}(h))$ (7)
This variation ensures the ultimate expert utilization [20] and has more compact model structure. However, in this paper, we mainly focus on the first type of CCoE framework and will left the variations as the future work.
4.2 Rule-based Gating
The inference procedure of CCoE firstly uses the shared layers to extract the upper level features of the inputs such as the syntax information, general linguistic patterns, etc. and then select one of experts for each query. To select the correct expert in CCoE, we need to know three information: the expert-domain query mapping, expert’s number layers and the position mask at expert layer because there could be several experts have the same number of layers. Another design requirement is to support flexibility. Since CCoE supports to continual train expert or add new expert, the CoE layers are dynamically changing. This requires the minimal change of our rule-based gating policy, if there is an update on one expert’s number of layers. We can rewrite the problem of rule-based gating as:
Consider a set of domain query prompts $Q=[q_{1},q_{2},\cdots,q_{n}]$ and a set of candidate experts $E$. The objective is to identify a sequence of executing paths $P=[P_{1},P_{2},\cdots,P_{n}]$, where $P_{i}$ contains the triplet $(\varphi_{i},l_{i},pos_{i})$, and $\varphi_{i}$ is $domain_{j}$-to-$expert_{i}$ mapping, $l_{i}$ is domain $expert_{i}$’ number of layers and $pos_{i}$ is the index of $expert_{i}$ at $L-l_{i}$ layer. For the simplicity, we use a rule-based task-expert matrix during the inference phrase. The matrix is a zero/one matrix $A_{m\times n}$, where $m$ is number of task and $n$ is the number of experts. At each $A_{i,j}$, it is a tuple structure that contains the number of layers and expert position at CoE layer. If new updated expert is coupled at the CoE layer, it will append to the last position to prevent the changes on Rule-based gating.
4.3 Expert Training
One of the main difficult in training MoE model is to balance the expert loads to prevent the case that tokens are always assigned to the same experts. Another problem is the knowledge redundancy. Knowledge sharing among experts means that different experts may have the repeated knowledge in their parameters. This leads to redundancy in expert parameters and prevent the expert specialization. To support specialized expert training, CCoE defines two types of operations: "push" and "pop". "push" means to add/update an expert in CCoE framework, and "pop" means deep copy/remove an expert in CCoE.
Figure 2: Example of training a new specialized expert with domain data and "push" into the CCoE model with 8 gpus
Figure 3: Example of continual training the specialized expert with updated domain data using "pop" and "push" operations with 8 gpus
In CCoE, the training can be divided into scenarios: the first scenario is to train a new expert LLM with a new domain data (Figure 1). It loads the new expert LLM $n$ into the gpu and feeds the domain tokens. After training, it creates a deep copy of the new expert ${n}^{\prime}$ in CCoE and transfers the weights to the ${n}^{\prime}$. After all, it will destroy the expert LLM $n$ and release the memory. Another scenario is to continually train an expert LLM (Figure 2) to allow the expert LLM to continually learn the domain knowledge. This can be summarized with 3 steps. Step 1. Creating a deep copy of expert $n$ at gpu 8 as expert ${n}^{\prime}$. Step 2. Feeding the domain tokens to continually train expert ${n}^{\prime}$. Step 3. Transferring the new weights of expert ${n}^{\prime}$ to CCoE expert $n$, destroy the expert ${n}^{\prime}$. The operations of "push" and "pop" support CCoE to add and continual train experts.
4.4 Expert LLM Architecture
The number of parameters of language model can varied based on the difficulty of the target tasks. Based on the scaling law, the model’s performance will improve when we increase the amount of model’s parameters and the training data. This enables model to continually increasing capabilities as we following the scaling law. However, in real-world scenario, it is a question that do we need to spend such amount of resources to train a giant language model for all downstream tasks, or a small model can achieve the comparable results with enough training tokens and training time. On the other side, recent research also realizes the deficit of training the large language model with large sizes such as the unstable training or token deficits. The training token should maintain the same ratio with sizes of large language models [21]. In most of case, dataset for downstream fine-tuning tasks may not have enough tokens, or small model can also achieve a comparable results as the large models and it is more economic in terms of training resources. Therefore, we defined a performance gains metrics to evaluate the improvement in expert performance with respect to the increase of expert parameters. This can be easily formulated as $\frac{-\Delta\varepsilon_{val}}{\left|{\Delta\theta}\right|}$. $\varepsilon_{val}$ and $\theta$ are the validation error and expert model’s number of parameters. If the model achieves the better performance, then there is a negative change in $\varepsilon_{val}$. If we increase the parameters but receive a small amount reduction in validation error, then value of performance gains would be small, and vice versa.
We also realize that most of the downstream tasks are related to natural language processing, especially for those applications that need interact with users. For example, building a Law-QA system is essentially a Question-Answering dialogue system with a Law-domain LLM. These specialized domain LLM share some common capabilities with pre-trained general LLMs in the perspective of natural language understanding, generation and inference. Therefore, it could be feasible to fine-tune several layers of the general LLM to inject domain knowledge rather than fine-tuning entire model.
The problem can be converted to determine the number of expert layers for different downstream domains. This demands a deep understanding about the underlying mechanism of LLMs. Recent researches propose the concept of Knowledge Neuron in LLMs [22], which are the neurons that express a relational fact. To be more elaborate, the mechanism underlying the ability of LLMs to recall facts in the training corpus. The naive way of interpreting KN thesis states that facts are stored into the MLP weights, in the manner of key-value memory [23]. However, recent work provides a more comprehensive understanding that looks beyond the MLP weights [24]. They believe that MLP weights operate on complex patterns that can be syntactically and semantically interpreted rather than only storing the factual knowledge. Since fine-tuning several layers of LLM can only modify portion of the MLP weights, it is questionable to prove the effectiveness of expert training in CCoE framework.
The study of interoperability in LLMs aims to understand the process of how LLMs encode tasks with varying complexities. Different LLM families with similar parameter scales demonstrate different behaviors of tackling down the same problem [25]. This reflects that the diverse way of different LLM families interpret and process complex information. Another interesting observations of interoperability has been made in terms of the model size. Large model from different families consistently achieve superior performance on difficult tasks than small models. Large LLMs are able to capture the difficult concepts at earlier layers. Therefore, we form the following research questions regarding to our CCoE expert training:
RQ1: What are the suitable sizes of experts that can balance the performance and model parameters for downstream tasks?
RQ2: Can fine-tuning several layers of LLM really make expert models achieve comparable performance in all downstream domains?
RQ3: Does the performance of expert is highly related to base model’s performance in terms of model size and model family?
Our training objective can be reformed as:
$argmin_{\mathcal{L}_{f_{\theta}}}=argmin_{\ell_{({\mathcal{L}_{i,t_{k}},\mathcal{L}_{j,t_{k}}}})}\forall i,j\in E,\forall t_{k}\in T\text{ }\textrm{s.t.}\text{ }min\mathcal{L}_{val}$ (8)
[…]
5 Experiments
5.1 Training Data
We collect 5 different domain datasets: Math, Code, Finance, law, Medical, and Text-to-SQL. Each domain focuses on enhancing model capabilities by injecting domain knowledge or providing reasoning training to experts. In our training corpus, we mainly used the publicly available and high quality datasets. The data composition is illustrated at Figure 3.
Math. Mathematical reasoning mainly relies on model’s inference capabilities and already lead the various approaches to improve the performance on math-related tasks. In our approach, we selected the recent math training dataset that has been shown to effectively boost model’s math performance. [3]. Besides, using the Chain-of-thought and RLHF technologies can further improve model’s math performance [26] [4]. We also constructed our training dataset by providing a step-by-step solution way rather than the vanilla format, and apply the dpo training to align with the preferences. In the our math dataset, we collected nearly 400K instructions mainly from MetaMathQA [^1] https://huggingface.co/datasets/meta-math/MetaMathQA.
Code. The code generation usually combines the execution results with structured code. These pairs are verified and filtered by code interpreter before being used to fine-tune large language models. To optimize the model’s reasoning capabilities, the data is usually formatted with COT prompting. We adopt the commonly used datasets from code instruction 128K alpaca  [^2] https://huggingface.co/datasets/iamtarun/code_instructions_120k_alpaca, code feedback [27] and MBPP [28]. Our merged dataset includes various programming languages and covers major programming topics. We organized the fine-tuning data into instruction-based format with over 200K number of pairs.
Text-to-SQL. The Text-to-SQL (NL2SQL) is a well-known challenge in NLP and database community. It targets to translate the natural language input into corresponding and valid Structured Query Language (SQL). It requires the LLM to understand the semantic purpose of natural language input, the target database schema and the grammar of SQL. Current proposed solution of this task includes SFT, or craftily designed prompt with paradigm of "prompt" learning or "in-context learning". We explore the SFT paradigm to empower LLM Text-to-SQL capability. We collect our dataset with total of 133K instruction-based pair and mainly from WikiSQL [29], Chase [30], and DuSQl [31].
Medical. The challenges of developing LLM-based medical applications are lack of factual inaccuracies, background knowledge for diagnoses, and insufficient grounding in real-world experience. There are many elaborately collected medical datasets that aim to optimize LLM capabilities in medical domain. We adapt the medical data from MedQA [32], MedicalGPT [33], and HuaTuo [34], and create a over 849K English and Chinese QA-paired medical dataset that mixes English and Chinese.
law. We mainly focus on the Chinese mainland law system because there is a possibility that some cases could have different decisions under the Common Law and Civil Law. We collected the instruction-based Civil Law dataset from some influential projects [35] [36]. Our law dataset includes a total of nearly 300K instructions-based pairs that involved with different tasks: decision prediction in a case, legal Q&A, judgment prediction, judicial reading comprehension, legal case classification, etc.
Figure 4: The data composition of our training corpus including the 5 different domains: Math, Code, Law, Medical, Text-to-SQL.
5.2 Data processing and training
The quality for open-sourced dataset is critical for training expert models. We firstly use Min-Hash and ’text-embedding-ada-002’ embedding model to remove highly similar questions. In downstream fine-tuning community, domain datasets often mix with pre-train dataset to avoid degradation of model’s general capabilities or they train the model with multiple-stages. However, in CCoE framework, each expert LLM fine-tuning only need to focus on learning the domain data. Our dataset preparation only needs to augment the domain data quality. We aim to overcome this through three dimensions: 1. comprehensive 2. complex 3. diverse. Based on our deliberately explorations on domain datasets, the Math dataset is mainly sampled from MetaMath [3]. It uses the bootstrapping approach to rephrase the math question with enough complexity, diversity and comphrensitiviy.
[…]
We also conducted the DPO training on Math, Code domains. To create the direct preference pairs, we ask the expert model to generate 5 outputs including the exact answers and reasoning steps with a sampling temperature of 0.8. We adopt the template form [39] to rank the answers and reasoning steps based on the dimensions of completeness, correctness, and readability. Comparing to directly rank the 5 model outputs, we decide to use the pairwise ranking to have a better stable ranking result. We conducted pairwise ranking among them and manually solved the conflicts if there are multiple outputs have the same ranking position or a loop in the ranking positions. We use the highest ranked output as the chosen answer and lowest ranked output as the rejected answer.
5.3 Model Settings and Evaluation
We use QWen1.5 [40] 7B and 13B parameters as our backbone models. To simplify our training process, we do not expand the vocabulary size and continue to train the word embedding for each domain. We conducted a comprehensive experiments on hyper-parameters $L$, the number of layer in each expert. We explored the model performance gains when we increase the $L$. In our context, we want to limit the layers of each expert to the upper bound $\left\lfloor\frac{L_{base}}{2}\right\rfloor$ because we want to control the overall CCoE model parameters under the constraints of computational resources. Many existing literature have already demonstrate that the deeper layers of LLM often acquire more difficult concepts and can be easily recognized for complex tasks [25] [41]. We start to fine-tune the layers starting from the last 4 layers and add 4 layers every time. We select the number of layers in each expert that gives the highest performance gains. The results of measuring performance gains are listed in Appendix.
Table 1: The comparison of CCoE model and existing LLMs on Math, Code, Text-to-SQL, Medical, and Law domains
GSM8K LawBench
7B model series 13-14B model series 6B-7B model series 13-14B model series
ToRA-Code 7B 72.6 ToRA-Code 13B 75.8 InternLM2-Chat-7B 48.53 Yulan-Chat-2-13B 34.5
MuggleMATH 7B 69.8 MuggleMATH 13B 74.0 Fuzi-Mingcha 28.78 ChatLaw-13B 32.63
MathCoder-CL-7B 67.8 MathCoder-CL-13B 72.6 Wisdom-Interrogatory 7B 27.74 Lawyer-LLaMA 23.02
MetaMath-Mistral-7B 77.7 KwaiYiiMath 13B 73.3 LexiLaw 6B 26.41 Ziya-LLaMA-13B 18.79
MetaMath 7B 66.4 MetaMath 13B 72.3 HanFei 7B 24.91 LLaMA-2-Chinese-13B 17.08
MathCoder-L-7B 64.2 MathCoder-L 3B 72.6 ChatGLM2-6B 22.94 BELLE-LLaMA-2 13B 30.79
WizardMath-7B-V1.0 54.9 WizardMath-13B-V1.0 63.9 Chinese-Alpaca-2-7B 19.22 LLaMA-2-13B 18.37
Baichuan2-7B 24.5 Baichuan2-13B 52.8 Baichuan2-7B 21.18 Vicuna-v1.3-13B 15.81
Qwen1.5-7B-chat 62.5 Qwen1.5-14B-chat 70.1 Qwen1.5-7B-chat 43.87 Qwen1.5-14B-chat 49.4
CCoE-7B Math 68.2 CCoE-14B Math 73.0 CCoE-7B Law 47.1 CCoE-14B Law 54.7
MBPP Human Eval
6-9B model series 13-14B model series 6-7B model series 13-16B model series
Code LLaMA Python 7B 47.6 Code LLaMA Python 13B 49 Code LLaMA Python 7B 38.4 Code LLaMA-Python-13B 43.3
CodeGemma-7B-IT 54.2 StarCoder 15.5B 53.2 CodeGen-Mono 6.1B 33.8 StarCoder 15.5B 33.6
DeepSeek-Coder-Base 6.7 60.6 CodeGeeX-13B 24.4 DeepSeek-Coder-Base 6.7B 44.7 CodeGeeX-13B 22.9
InCoder 6.7B + CodeT 34.4 WizardCoder 15B 51.8 OpenChat-3.5 7B 55.5 InstructCodeT5+ 16B 35.0
Gemma 2 - 9B 52.4 CodeGeeX-13B 24.4 CodeT5+6B 28.0 PanGu-Coder2 61.6
CodeGemma-7B 56.2 CodeGen 16B + Coder-Reviewer 46.2 CodeGemma-7B 44.5 WizardCoder-15B-V1.0 57.3
Code Llama - Instruct 7B 54.9 code-cushman-001 12B 55.4 CodeGeeX2-6B 35.9 CodeT5+ 16B 30.9
Baichuan2-7B 24.2 Baichuan2-13B 52.8 Baichuan2-7B 18.29 Baichuan2-13B-Base 17.07
Qwen1.5-7B-chat 37.4 Qwen1.5-14B-chat 44.0 Qwen1.5-7B-chat 36.0 Qwen1.5-14B-chat 37.8
CCoE-7B Code 45.5 CCoE-14B Code 48.8 CCoE-7B Code 40.8 CCoE-14B Code 43.7
MedQA Text-to-SQL
Med-PaLM 2 79.2 CoH + GPT4 83.9
MeerKat-7B 74.3 CoH + GPT4 + instructions 82.3
CCoE-14B Med 73.5 C3 + ChatGPT 82.3
Meditron-70B 70.2 RESDSQL-3B + NatSQL 79.8
Flan-PaLM 540B 67.6 Flan-PaLM 540B 67.6
CCoE-7B Med 65.4 DIN-SQL + CodeX 78.2
Qwen1.5-14B-chat 62.3 BRIDGE v2 + BERT(ensemble) 68.3
Meditron-70B 61.5 BRIDGE + BERT 59.9
Qwen1.5-7B-chat 57.8 CoH + CCoE-14B SQL 56.1
BioMedGPT-10B 50.4 CoH + CCoE-7B SQL 41.8
PubMedGPT 2.7B 50.3 Qwen1.5 14B chat (4-shot) 34.5
BioBERT large 36.7 Qwen1.5 7B chat (4-shot) 29.9
Table 1 shows the testing accuracy of our 5 domain experts on GSM8K [^3] https://paperswithcode.com/sota/arithmetic-reasoning-on-gsm8k, MBPP [^4] https://paperswithcode.com/sota/code-generation-on-mbpp, Human Eval [^5] https://paperswithcode.com/sota/code-generation-on-humaneval, LawBench  [^6] https://lawbench.opencompass.org.cn/leaderboard, MedQA [^7] https://paperswithcode.com/sota/question-answering-on-medqa-usmle, and Spider [^8] https://yale-lily.github.io/spider benchmarks. Our CCoE framework is able to achieve the comparable performance and have the greatly outperformed the baseline model Qwen1.5 7B chat and 14B chat [^9] https://qwenlm.github.io/blog/qwen1.5/. Our results illustrate that there are nearly 10%-20% improvement on the baseline models in 5 domains. This indicates that the CCoE framework can easily boost the base model performance in different downstream domains. However, expert in Math has the less improvements after fine-tuning (CCoE 14B Math). We think the base model already achieves a comparably high test results on the benchmark. Our current training strategies can only play limited contributions.
To answer the RQ2, Law and medical domains need to have more fine-tuning layers to delivery better test results. We think this is due to that they have less portion in pre-training corpus and some of their knowledge are contemporary. To revise the contemporary knowledge requires to correct the knowledge neurons. Having more fine-tuning layers allows us to modify the knowledge neurons in the lower layers. However, Math and Code may have larger portions during the pre-training and already have enough knowledge. Our current fine-tuning emphasizes on improving model’s reasoning capability through CoT, question bootstrapping.[3], and self-verification [6].
The Text-to-SQL requires the LLM to understand the SQL syntax, database schema, and the capability of LLM is also very essential. We observed that the Qwen1.5 base model can give the some exact and partial correct results with few-shot prompt. However, they fail to generate many correct SQLs, especially for those hard-level questions. Besides, even though we have conducted the fine-tuning, there is still a performance gap between the models on Spider Leader-board. Comparing to the base model, there is a big improvement. We only conducted limited study on the effect of prompting in Text-to-SQL. In the future, we plan to continue to study the In-Context Learning and calibration of hints to improve the Text-to-SQL performance.
To answer the RQ 3, We select the open-sourced LLaMA2 7B chat model [42] as our comparison model. LLaMA2 7B has the same model scale as the QWen1.5 7B model. To exclude other confounding variables, we train the LLaMA2 7B chat model with same number of expert layers and same dataset. From table 4, we can see that LLaMA model is able to receive a higher performance boost than Qwen1.5 base model. We speculate that a comparatively low performance base model can easily to be enhanced through the downstream fine-tuning and data augmentation strategies. However, LLaMA-based experts still receive low results in 5 domain evaluations. We think this is because expert training requires to utilize the lower layers of the base model to extract the language or even semantic features, which are still decided by the base model’s capabilities. We can conclude that the CCoE framework can boost the performance of base model in different downstream domains regardless of model’s family but expert’s performance is constraint on its base model’s capability. We also realize the scaling law applies to the expert training. It is difficult to continue to improve expert’s performance by increasing the number of training data without changing the expert size.
Table 2: The comparison between LLaMA2 7B chat and Qwen1.5 7B chat as the base model of CCoE framework
Domain CCoE-LLaMA2 7B Chat CCoE-Qwen1.5 7B chat
MBPP 32.3(+11.5) 45.5(+8.1)
Human Eval 21.4 (+8.6) 40.8 (+4.8 )
GSM8K 36.6(+19.9) 68.2(+5.7)
Law 14.5(+7.35) 47.1(+3.3)
Med 30.4(+8.0) 65.4(+7.6)
Text-2-SQL 30.7(+16.2) 41.8 (+11.9)
5.4 Limitations
Our proposed CCoE framework depends on adding expert layers into the base model to form different experts. In this work, we only conducted preliminary study on injecting knowledge to the last several layers. This may not allow the model to give the best performance. More explorations need to be done on training strategies and data compositions, or more comprehensive experiments are needed to identify the best expert layers, which could be upper, middle or distributed among the base model. We could utilize the similar approach as locating the Knowledge Neurons to find out the most appropriate combination of expert layers that enable the expert to achieve the best performance. Furthermore, we also realize that there is the performance upper bound under the constraint of only fine-tuning the half of the base layers. However, we believe that it is a trade-off between the model’s performance and model’s size.
6 Summary and Future Work
In this paper, we propose the CCoE framework that aims to boost the base model’s overall performance in all perspectives through utilizing the idea of collaborating multiple of experts (CoE) LLM layers. All of the expert LLM layers are coupled at the base model and becomes the expert sub-network. We conducted the experiments on Math, Code, Law, Medical and Text-to-SQL domains. Our results indicate that there are at nearly 10%-20% improvements on these 5 downstream domains. This proves that CCoE framework is able to boost the overall performance of the base model with the advantages of low-cost, high-interpretability and efficiency. Moreover, CCoE also has the high extensibility through scaling up the number of expert layers to fulfill more open-ended tasks. However, the current routing of the expert layers depends on the API endpoints trace. In the future, We plan to use the LLM to help us route the prompt to the correct expert sub-network. We can extend the prompt’s content to ask the base model to classify the prompt’s task or domain during the first generation and then run the second inference to route the real prompt to the correct expert sub-network. We understand this could result in the high Time to First Token (TTFT), but there is still great potential of the CCoE network for us to explore.
Appendix A Appendix
Table 3: ChatGPT Prompting Template for pairwise ranking expert model outputs
Instruction:
\$$\$\{instruction\}$
Input:
\$$\$\{input\}$
Candidate A:
\$$\$\{candidateA\}$
template Candidate B:
\$$\$\{candidateB\}$
Given the instruction and input above, please compare the two
candidates based on output’s completeness, correctness and readability.
Each candidate should include the final answer, a step-by-step solution and test cases.
You only have 4 choices to output:
If you think A is better, please output: 1. A is better
if you think B is better, please output: 2. B is better
If you think both are good enough and satisfy the three requirements, please output: 3. Same good
If you think both are bad and do not follow instruction or satisfy three requirements, please output: 4. Same bad
Please strictly output the above 4 choices and do not output anything else.
Output your choice below:
1. A is better
2. B is better
Comparison Option 3. Same good
4. Same bed
Table 4: Example prompt for evaluating the baseline of Text-to-SQL
Task: Reference the following examples. Given the database schema and
question, answer the sqlite SQL query with no explanations.
Example 1: Given the following database schema:
[…]
Example 2: Given the following database schema:
[…]
$\cdots$
Given the following database schema:
[…]
Answer: $\cdots$
Table 5: Example prompt for fine-tuning Text-to-SQL expert
Below is an instruction that describes a task, paired
with an input that provides further context. Write a
response and with no explanation that appropriately completes the request
### Instruction:
Write a sql to answer the question: "How many clubs are there?"
### Input:
club(club id, name, manager, captain, manufacturer, sponsor)
player(player id, name, country, earnings, events number, wins count, club id)
### Response:
$\cdots$
Figure 5: The performance gains of different CCoE expert layers in 5 domains after fine-tuning